OpenStudy (kaitlyn_nicole):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2.

OpenStudy (123ab456c):

y=−x+4 and gx=−2x−3

OpenStudy (kaitlyn_nicole):

yeah? @123AB456C

OpenStudy (123ab456c):

YEP!:) go to mathpapa and go to the algebra calculator!:):)

OpenStudy (kaitlyn_nicole):

i wanna learn how to do it...

jhonyy9 (jhonyy9):

like a first step calculi the g(2)=? and after this just calculi the f(g(2)) = ?

OpenStudy (kayders1997):

Hi there

OpenStudy (kayders1997):

So when you get a f(g(x)) that means you are putting the function g into f, so always work your way from inside to outside

jhonyy9 (jhonyy9):

yes in this case sure - ok. ?

OpenStudy (kayders1997):

So start with you g function and put it into f like this... -(-2x-3)+4 so your just putting in g for the x

OpenStudy (kayders1997):

Than distribute and put 2 in for x and than add it all up and you should get answer

OpenStudy (kayders1997):

Do you understand?

jhonyy9 (jhonyy9):

no - i have said first you calculi the value of g(2) = ? and after that you just calculi the f(g(2)) so g(2) you get a number - ok ?

OpenStudy (kayders1997):

Whats your answer?

OpenStudy (kayders1997):

@jhonyy9

jhonyy9 (jhonyy9):

nicole can you calculi the g(2) = ?

OpenStudy (kayders1997):

I think you get the same answer either way the way I did it I got 11

OpenStudy (kayders1997):

Your way is a lot easier though @jhonyy9 :)

jhonyy9 (jhonyy9):

g(2) = -7 so this mean that f(-7) = -(-7)+4 = ?

jhonyy9 (jhonyy9):

yes @kayders1997 ?

OpenStudy (kayders1997):

Yes that what I got

OpenStudy (kayders1997):

*thats

jhonyy9 (jhonyy9):

Nicole - do you understand it now ?

OpenStudy (kaitlyn_nicole):

what the heck

OpenStudy (kaitlyn_nicole):

no not at all tbh

jhonyy9 (jhonyy9):

Nicole so than you dont understand it please ask me what you wan please ?

OpenStudy (kayders1997):

What don't you get? What part?

OpenStudy (kayders1997):

It's okay to not understand it right away, are you looking at her work my work or both?